### v3

# A Volumetric Approach to Privacy of Dynamical Systems

Abstract
Information-theoretic metrics, such as mutual information, have been widely used to evaluate privacy leakage in dynamic systems. However, these approaches are typically limited to stochastic systems and face computational challenges. In this paper, we introduce a novel volumetric framework for analyzing privacy in systems affected by unknown but bounded noise. Our model considers a dynamic system comprising public and private states, where an observation set of the public state is released. An adversary utilizes the observed public state to infer an uncertainty set of the private state, referred to as the inference attack. We define the evolution dynamics of these inference attacks and quantify the privacy level of the private state using the volume of its uncertainty sets. We then develop an approximate computation method leveraging interval analysis to compute the private state set. We investigate the properties of the proposed volumetric privacy measure and demonstrate that it is bounded by the information gain derived from the observation set. Furthermore, we propose an optimization approach to designing privacy filter using randomization and linear programming based on the proposed privacy measure. The effectiveness of the optimal privacy filter design is evaluated through a production-inventory case study, illustrating its robustness against inference attacks and its superiority compared to a truncated Gaussian mechanism.
keywords:
Volumetric privacy measure; privacy protection; interval analysis; truncated Gaussian mechanism.
thanks: The work was partially supported by the Research Grants Council of Hong Kong under Project CityU 21208921, a grant from Chow Sang Sang Group Research Fund sponsored by Chow Sang Sang Holdings International Limited.
1 Introduction
1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information. Specifically, correlations between shared metrics and underlying operational parameters can be exploited by adversaries to develop competitive and malicious strageties. This challenge highlights the critical need for methodologies that preserve data utility while ensuring rigorous privacy protection for dynamic systems.
Information-theoretic metrics, including mutual information, directed information, and conditional entropy, have been extended to quantify privacy leakage in dynamic systems. Despite their utility, these metrics have notable limitations. For instance, as highlighted in [8, 30], mutual information evaluates privacy leakage in an average sense, which can overlook privacy breaches for infrequent realizations of private variables. The computational complexity of these measures is another challenge, as they require evaluating leakage for all possible realizations of random variables [25, 32, 11]. Most existing information-theoretic approaches are focused on stochastic systems and typically assume complete knowledge of the probability distributions of system states. However, this assumption does not hold in systems that are influenced by unknown but bounded (UBB) noise.
To address these challenges, this paper introduces a volumetric privacy measure based on the uncertainty set of the private variable. The proposed approach is applicable to both deterministic and stochastic systems and eliminates the need for prior knowledge of probability distributions.
1.2 Contributions
This paper addresses the privacy protection problem for dynamic systems with UBB noise, as illustrated in Fig. 1. In this framework, the system state is partitioned into two categories: the public state $X_{k}$ and the private state $Y_{k}$, both of which belong to a known bounded set. The observation of the system states is represented by the set $\mathcal{M}_{k|k}^{x}$ which contains the actual public state $X_{k}=x_{k}$. An untrusted third party, i.e., the adversary, uses $\mathcal{M}_{k|k}^{x}$ to infer the private state by constructing an uncertainty set $\mathcal{Y}_{k|k}$, which is referred to as the inference attack.
Figure 1: The inference attack.
This system setup is inspired by the production-inventory problem discussed in Section 2. In such scenarios, inventory information can aid distributors in optimizing sales. However, it also enables adversaries to infer sensitive information, such as the production rate, which reveals a company’s private data, including production efficiency, strategies, and supply chain dynamics.
The contributions of this paper are summarized as follows:
  • Definition of Inference Attack: We define the adversary’s inference attack using set operations.
  • Volumetric Privacy Measure: A novel privacy measure is proposed based on the volume of the uncertainty set for the private state.
  • Privacy Level Computation: We develop a computational method utilizing interval analysis to evaluate the privacy level.
  • Optimal Privacy Filter Design: An optimization-based approach is introduced to design privacy filters, based on randomization and linear programming.
Additionally, we validate our approach using the production-inventory example. The results demonstrate that the adversary can infer sensitive production rates, through inventory data. However, this leakage can be significantly mitigated under the proposed optimal privacy filter design. Also, compared with the truncated Gaussian mechanism, our privacy filter design achieves lower data distortion while maintaining the same privacy level.
1.3 Related work
Differential privacy was initially defined to hide the presence of an individual or a specific record in the responses of queries for static datasets. The authors in [15] extended the differential privacy to dynamic systems, and proposed the differentially private Kalman filter for linear systems. In [5], the authors proposed a general architecture for private filtering by approximating the desire filter in a differential private way, and studied the application of differential privacy in distributed control and optimization. In [21], the authors studied the differential privacy preserving average consensus problem to protect the initial state in distributed control by adding and abstracting random noises. The authors in [12] connected the input observability with differential privacy and proposed the privacy preserving controller design method. The authors in [33] studied the minimal amount noises added to multi-agent systems for differential privacy based on the minimal observability subspace. They proposed node-based and edge-based privacy-preserving mechanisms based on the optimization of the minimum added noise, and discussed the trade-offs between utility and privacy. In [34], the authors proposed several cost-friendly differential privacy-preserving schemes by randomly adjusting the charge-discharge rate of a batter based on a modified Laplace distribution.
Information-theoretic protection methods use conditional entropy, mutual information, etc, from information theory, to measure privacy leakage. In [22], the authors studied the privacy filter design for a Markov chain to hide the private state when transmitting the monitored state to a receiver, which was approximately solved with a greedy algorithm based on the convex optimization. The authors in [3] designed a private filter to protect the state of a hidden Markov chain via minimizing the mutual information with the constraint of the observation utility. In [31, 24], the authors studied the privacy-aware filter design for linear systems, using the directed information between the system state and the output of the filter as the privacy measure. It was proved that the optimal filter can be realized with a Kalman filter and an additive Gaussian mechanism.
The authors in [32] studied the structural properties of the privacy-aware state estimation problem based on the dynamic programming decomposition. They approximated the privacy leakage with variational techniques and solved the estimation problem via policy gradient approaches. In [23], the conditional entropy of the private state given measurements and controls was used as a penalty term in the partially observable Markov decision process (POMDP) optimization, to avoid the adversary accurately estimate the system state. They showed that the state obfuscation problem is a standard POMDP optimization problem in which the cost-to-go function is concave in the belief state. In [16], the authors considered to protect the privacy of the electricity demand of a household to avoid the adversary infers the indoor activity. They utilized the rechargeable battery to hide the actual electricity demand and obtained the optimal charging policy via minimizing the privacy leakage measured by mutual information. However, the classical information-theoretic metrics for privacy might have some limitations when facing one-try attack as discussed in [30], which motivates the development of $R\acute{e}nyi$ min-entropy [30], maximal leakage [9] and pointwise maximal leakage [28], etc.
In most existing privacy protection literature for dynamic systems, it is assumed that the system are disturbed with unbounded noise that has probabilistic properties, e.g., Gaussian distribution. However, there are also dynamic systems are driven with UBB noise without probability assumptions. The privacy concept for these systems are not well developed. The authors in [6, 7] designed differentially private set-based estimator to protect the privacy of measurements based on a truncated noise distribution. In [13], the authors proposed the guaranteed privacy extent from differential privacy to hide the measurement sensor identities, and provided the optimization method for the $\mathcal{H}_{\infty}$ optimal privacy preserving interval observer design. Furthermore, in [26, 27], authors studied the state opacity problems where the discrete state space is divided into disjoint sets of secrete and non-secrete states, and the system outputs from secrete and non-secrete sets are indistinguishable. The authors in [19] proposed $\delta-$approximate initial-state opacity as a secrete notion for discrete-time systems with continuous states, and developed verification approaches based on Barrier certificates.
The proposed volumetric privacy measure in this work is inspired by set-membership estimation methods for systems with UBB noise [20]. Most set-membership methods represent states as bounded geometric sets, such as intervals [10], zonotopes [14], or ellipses [4], where the volumes of these sets correspond to estimation uncertainty. Building on these methods, we consider the adversary’s ability to infer the private state set based on observations of the public state. Consequently, a larger volume of the uncertainty set implies a broader range of potential private states that could be consistent with the observed public state set. This motivates our use of the volume of the private state set as a privacy measure.
This work differs from prior studies, such as [6], [7], and [13], which focus on ensuring the differential privacy of system states. In contrast, we address the problem of protecting private states while publicly releasing non-sensitive states, as illustrated in Fig. 1. Our approach uses a volumetric privacy measure, which also distinguishes it from state opacity problems [26], [27], and [19], where the state space is partitioned into secret and non-secret sets.
Furthermore, state opacity problems focus on guaranteeing privacy by ensuring that at least one non-secret state produces an observation indistinguishable from the secret state. In these frameworks, enlarging the volume of the non-secret state set does not enhance privacy, as privacy is determined solely by the existence of indistinguishable outputs, not by the size of the state set. Differently, within our framework, a larger private state set is preferred, as it increases the adversary’s uncertainty in inferring the private state.
1.4 Outline
The rest of the paper is organized as follows. Section 2 introduces the system model and the inference attack. Section 3 defines the volumetric privacy measure, provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure. Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level. Section 5 presents numerical results, followed by the conclusions in Section 6.
1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. We use the vector $\left[\begin{array}[]{c}\underline{X}\\ \overline{X}\\ \end{array}\right]$ to describe the interval $\mathcal{X}$, i.e., $\mathcal{X}=\left\{X|\underline{X}\leqslant X\leqslant\overline{X}\right\}$, where $\underline{X}$ and $\overline{X}$ are the lower and upper endpoints, respectively. The interval $\mathcal{X}$ can also be represented with
$\mathcal{X}=\left\{c^{x}+\mathrm{diag}\left(p^{x}\right)\alpha:\alpha\in\mathcal{R}^{n_{x}},\left|\alpha\right|_{\infty}\leqslant 1\right\},$
where $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and $p^{x}=\frac{\overline{X}-\underline{X}}{2}$ are the center point and radius of the interval. Besides, given the block matrix $A=\left[A_{1},A_{2}\right]$, the multiplication between the matrix $A$ and the interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$. Also, if $\mathcal{X}$ and $\mathcal{Y}$ are intervals, then we use $\mathcal{X}+\mathcal{Y}$ to denote the interval $\left[\begin{array}[]{c}\underline{X}+\underline{Y}\\ \overline{X}+\overline{Y}\\ \end{array}\right]$, and use $\mathcal{X}-\mathcal{Y}$ to represent the interval $\left[\begin{array}[]{c}\underline{X}-\overline{Y}\\ \overline{X}-\underline{Y}\\ \end{array}\right]$. As for the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the new set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z+R|X\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\times n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$.
2 System Model and Inference Attack
2.1 System Model
We consider the following stable system model $\mathbf{G_{1}}$,
$\mathbf{G_{1}}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ \end{array}\right.,$ (3)
where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the UBB noises. Also, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively. When $A_{3}$ is zero, the model $\mathbf{G_{1}}$ can be regarded as a dynamic system with the public state $X_{k}$ and the private input $Y_{k}$.
For analysis convenience, we assume that $\mathcal{X}_{0|-1}$, $\mathcal{Y}_{0|-1}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$ are intervals. For other types of bounded sets, such as zonotopes [14] and ellipses [4], the analysis is left for future investigation. Note that we do not assume the prior knowledge of probability distributions of the process noises, e.g., Gaussian distribution, which is different from most existing literature in privacy protection for dynamic systems. Furthermore, we assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
2.2 Motivating Example
Consider the problem of privacy leakage related to the production rate in inventory control systems. To enhance sales, companies often need to share inventory levels with distributors. However, inventory levels are closely correlated with the production rate, which serves as a key indicator of the factory’s production strategy, operational efficiency, and supply chain dynamics. Therefore, the production rate constitutes highly sensitive information. If accessed by an adversary, this data could be exploited to develop competitive strategies, potentially compromising the company’s market position and operational integrity.
The inventory of production can be modeled as follows [17, 29],
$X_{k}=X_{k-1}+Y_{k-1}-W_{k}^{x},$ (4)
where $X_{k}$ is the inventory, $Y_{k}$ is the production rate, and $W_{k}^{x}$ is the uncertain demand. The production rate is usually chosen to maintain the inventory around a desired level, which can be modeled as follows,
$Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+W_{k}^{y},$ (5)
where the uncertain process noise $W_{k}^{y}$ is due to the adversary’s limited knowledge about the company’s production capability.
When the uncertainty of demand and production rate is small, e.g., $W_{k}^{x}=D^{x}$ and $W_{k}^{y}=D^{y}$ are constant, then the adversary can approximately reconstruct the production rate with the inventory level via
$Y_{k-1}\approx X_{k}-X_{k-1}+D^{x},$ (6)
and
$Y_{k}\approx A_{3}X_{k-1}+A_{4}Y_{k-1}+D^{y},$ (7)
which causes privacy leakage and increases potential risks of business competition.
However, in practical scenarios, the uncertainty associated with demand and production capability is bounded but cannot be ignored. Inferring the private production rate, therefore, becomes a complex but still feasible task. Due to the presence of unknown noise terms that belong to bounded sets without probabilistic assumptions, multiple private states may correspond to the same public state with equal probability, leading to the formation of an uncertainty set. To address this challenge, we define the adversary’s inference attack on the private state using set operations in the following sections.
2.3 Inference Attack
We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$. Also, every element of $\mathcal{M}_{k|k}^{x}$ is considered a potential candidate for the actual public state. Notably, if the public state is directly transmitted to the adversary without any processing, the public state set $\mathcal{M}_{k|k}^{x}$ contains only the actual public state.
Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set. This process is referred to as the inference attack. We next define the inference attack recursively.
At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of public state can be predicted based on the system model (3), i.e.,
$\mathcal{X}_{k|k-1}=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x}.$ (8)
Similarly, the set of private state can be predicted via
$\mathcal{Y}_{k|k-1}=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (9)
After receiving the observation set of the public state $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$, the adversary extracts new information from $\mathcal{M}^{x}_{k|k}$ and updates the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via the following steps,
$\mathcal{M}_{k-1|k}^{x}=$$\left\{M^{x}_{k-1|k}|X_{k|k}=A_{1}M^{x}_{k-1|k}+A_{2}Y_{k-1|k-1}+B_{1}W_{k}^{x},\right.$
$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall Y_{k-1|k-1}\in\mathcal{Y}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (10)
$\mathcal{M}_{k-1|k}^{y}=$$\left\{M^{y}_{k-1|k}|X_{k|k}=A_{1}X_{k-1|k}+A_{2}M^{y}_{k-1|k}+B_{1}W_{k}^{x},\right.$
$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall X_{k-1|k-1}\in\mathcal{X}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (11)
$\mathcal{X}_{k-1|k}=\mathcal{M}_{k-1|k}^{x}\cap\mathcal{X}_{k-1|k-1},$ (12)
$\mathcal{Y}_{k-1|k}=\mathcal{M}_{k-1|k}^{y}\cap\mathcal{Y}_{k-1|k-1},$ (13)
where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (2.3) and (2.3), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (12) and (13).
According to the system dynamics (3), the adversary estimates the uncertainty set of $Y_{k}$ via the following forward inference,
$\mathcal{Y}_{k|k}=A_{3}\mathcal{X}_{k-1|k}\oplus A_{4}\mathcal{Y}_{k-1|k}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (14)
Finally, the public state set can be further calibrated via intersection,
$\mathcal{X}_{k|k}=\mathcal{M}^{x}_{k|k}\cap\mathcal{M}^{x}_{k|k-1},$ (15)
where
$\mathcal{M}^{x}_{k|k-1}=A_{1}\mathcal{X}_{k-1|k}\oplus A_{2}\mathcal{Y}_{k-1|k}\oplus B_{1}\mathcal{W}_{k}^{x},$ (16)
is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (2.3)-(13), and the forward inference (14)-(16). The backward calibration (2.3)-(13) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
Proposition 1.
For any $k\geqslant 1$, the adversary’s uncertainty private state set (14) is a subset of its prediction set (9). Also, given the uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ containing the actual system states $X_{k-1}=x_{k-1}$ and $Y_{k-1}=y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the actual public state $X_{k}=x_{k}$, then the inference result $\mathcal{Y}_{k|k}$ also contains the actual private state $Y_{k}=y_{k}$.
Proof 2.1.
The proof is omitted due to the simplicity.
According to Proposition 1, the inference attack results in a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$, meaning that there are fewer possible private states corresponds to the same observation set $\mathcal{M}_{k}^{x}$. In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain actual private state. In conclusion, the adversary can reduce its uncertainty of the private state via the inference attack.
2.4 Privacy Measure and Defense
The aforementioned inference attack relies on set operations, and its outcome is the uncertainty set of the private state, denoted as $\mathcal{Y}_{k|k}$. The first key question addressed in this paper is how to evaluate the privacy level of the private state using the uncertainty set $\mathcal{Y}_{k|k}$. Also, set operations, such as the addition in (2.3) and the intersection in (12), involve all possible combinations of elements from different collections. These operations are computationally challenging due to the continuous nature of the state space. Consequently, it is essential to provide efficient computational tools for privacy level evaluation. Moreover, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage.
To address these challenges, we propose a volumetric privacy measure and its computation approach in Section 3. Furthermore, we present an optimal privacy filter design to reduce privacy leakage in Section 4.
3 Volumetric Privacy Measure
In this section, we define the volume of uncertainty sets as a quantitative measure of the privacy level. We then introduce an interval-based approach to approximate the computation of state uncertainty sets and evaluate the corresponding privacy levels. Additionally, we analyze the properties of the volumetric measure, showing that both the privacy level and the uncertainty reduction are bounded by the new information derived from the observation set.
3.1 Privacy and Utility Measures
As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$. However, due to the continuous nature of the state space, the number of elements in such sets is uncountable. To address this, we propose using the volume of the uncertainty set as a quantitative measure of privacy, defined as follows,
$P_{k}\left(\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$ (17)
where $\mathrm{Vol}\left(\cdot\right)\geqslant 0$ is a Lebesgue measure. When $\mathcal{Y}_{k|k}$ is an interval, we can compute its volume via $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\sum_{i=1}^{n}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$, where $\overline{Y}_{k|k}$ and $\underline{Y}_{k|k}$ are the upper and lower endpoints of the interval $\mathcal{Y}_{k|k}$, respectively.
Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, specially, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero. On the other hand, if the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is large, then the range of private state will be large, which increases the difficulties of selecting the correct value as the estimation result. Therefore, the volume of $\mathcal{Y}_{k|k}$ describes the amount of the adversary’s inference uncertainty about the private state as well as the privacy level. We can increase the value of $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ to increase the privacy level to protect the system from inference attack.
Besides, since the uncertainty set $\mathcal{Y}_{k|k}$ is a subset of its predicted version $\mathcal{Y}_{k|k-1}$, we have
$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right),$ (18)
meaning that the adversary’s inference uncertainty about the private state is bounded by its prior knowledge $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$. In other words, the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ cannot exceed $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$.
Furthermore, we propose to use the following measure to measure the utility of public state set $\mathcal{X}_{k|k}$,
$U_{k}\left(\mathcal{X}_{k|k}\right)=\frac{1}{\mathrm{Vol}\left(\mathcal{X}_{k|k}\right)},$ (19)
since a larger $\mathcal{X}_{k|k}$ provides more possible values of the public state to the receiver, i.e., increases the estimation error. Consequently, a larger value of $U_{k}\left(\mathcal{X}_{k|k}\right)$ indicates a higher data utility of the public state set.
3.2 Inference Attack Approximation
In this subsection, we investigate the inference approximation for multi-dimensional systems using the tightest interval method [1]. This approach simplifies the computation of the privacy measure and provides the foundation for designing the optimal privacy filter as discussed in Section 4.
Lemma 2.
The tightest recursive inference interval from (2.3) to (13) can be computed via
$\mathcal{M}_{k-1|k}^{x}=$ $\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
$+\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (20)
$\mathcal{M}_{k-1|k}^{y}=$ $\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{Y}_{k-1|k-1}$
$+\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (21)
$\mathcal{X}_{k-1|k}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{x},\underline{X}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{x},\overline{X}_{k-1|k-1}\right\}\\ \end{array}\right],$ (24)
$\mathcal{Y}_{k-1|k}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{y},\underline{Y}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{y},\overline{Y}_{k-1|k-1}\right\}\\ \end{array}\right],$ (27)
$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!+\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (28)
$\mathcal{X}_{k|k}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k|k}^{x},\underline{M}_{k|k-1}^{x}\right\}\\ \min\left\{\overline{M}_{k|k}^{x},\overline{M}_{k|k-1}^{x}\right\}\\ \end{array}\right],$ (31)
$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}=\Psi\left(A_{3}\right)\mathcal{X}_{k-1|k}+\Psi\left(A_{4}\right)\mathcal{Y}_{k-1|k}+\Psi\left(B_{2}\right)\mathcal{W}_{k}^{y},$ (32)
with
$\Psi\left(\star\right)=\left[\begin{matrix}\frac{\star+\left|\star\right|}{2}&\frac{\star-\left|\star\right|}{2}\\ \frac{\star-\left|\star\right|}{2}&\frac{\star+\left|\star\right|}{2}\\ \end{matrix}\right].$
Also, the tightest prior inference set of $Y_{k}$ is
$\!\!\!\!\!\!\!\mathcal{Y}_{k|k-1}\!=\!\Psi\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y},$ (33)
if $k\geqslant 1$. If $k=0$, then $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$ and
$\mathcal{X}_{0|0}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{0|0}^{x},\underline{X}_{0|-1}\right\}\\ \min\left\{\overline{M}_{0|0}^{x},\overline{X}_{0|-1}\right\}\\ \end{array}\right].$ (36)
Proof 3.1.
See Appendix A.
According to Lemma 2, the inference attack can be approximately computed using intervals. It can be verified that the intervals computed in Lemma 2 are tight when $X_{k}$ and $Y_{k}$ are scalar variables.
3.3 Properties of the Inference Attack
The inference attack exhibits certain properties. For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the process noise and the observation set, as stated below.
Lemma 3.
For any $k\geqslant 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
$p_{k|k}^{y}\leqslant$ $\left(\left|A_{3}\right|+\left|A_{4}\right|\left|A_{2}^{-1}\right|+\left|A_{4}\right|\left|A_{2}^{-1}A_{1}\right|\right)\overline{p}^{x}$
$+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p^{w,x}_{k}+\left|B_{2}\right|p^{w,y}_{k},$ (37)
where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively.
Proof 3.2.
See Appendix B.
Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of $\overline{p}^{x}$. As a result, if the value of elements in $\overline{p}^{x}$ are small, i.e., if the observation set $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level is low, and the adversary experiences less uncertainty after performing the inference attack.
Furthermore, by comparing the predicted and posterior uncertainty sets, as given by (9) and (14), the amount of uncertainty reduction can be quantified by $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, where
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\diagdown\mathcal{Y}_{k|k}.$
Proposition 4.
If $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, then the volume of $\Delta\mathcal{Y}_{k|k}$ is
$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$
Proof 3.3.
Since $\mathcal{Y}_{k|k}$ is the subset of $\mathcal{Y}_{k|k-1}$, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is equal to $\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$, which can be computed as follows,
$\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$
$=$ $\sum_{i=1}^{n_{y}}\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$
$=$ $\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k-1}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)}-\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$
$=$ $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$
where $\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$ is the length of i-th interval of $\Delta\mathcal{Y}_{k|k}$.
According to Proposition 4, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
Theorem 5.
The amount of uncertainty reduction at $k$ is
$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}+\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (38)
satisfying
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}=\mathcal{X}_{k-1|k-1}-\mathcal{X}_{k-1|k}$ and $\Delta\mathcal{Y}_{k-1|k}=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$, $\left\|A\right\|\frac{\Delta}{=}\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$ is the absolute value of $A$.
Proof 3.4.
See Appendix C.
According to Theorem 5, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$. Furthermore, with Proposition 4 and Theorem 5, we have the following lemma to bound the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$.
Lemma 6.
The privacy level can be bounded with the following inequalities,
$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
$\leqslant\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
Therefore, on the one hand, one can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$. On the other hand, when the private level is high, we also avoid the adversary to update its central estimation $c_{k|k}^{y}$ due to the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
4 Privacy Filter Design Problem Using the Volumetric Privacy measure
In this section, we address the privacy filter design problem, where the filter outputs an appropriate observation set that balances the trade-off between the data utility of the public state and the privacy protection of the private state.
4.1 The Structure of Privacy Filter
We begin by defining the decision domain of the privacy filter as follows. At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the tightest interval of $X_{k}$ can be computed via Lemma 9, i.e.,
$\!\!\mathcal{X}_{k|k-1}\!=\!\Psi\!\left(\!A_{1}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\!\left(\!A_{2}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x}\!,$
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subset\mathcal{X}_{k|k-1}$. To maintain the high data utility, the observation set has to satisfy the following constraint,
$\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}.$
We next consider the privacy filter drawn in Fig.2, which first randomly generates a set $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$ that contains the actual public state $X_{k}=x_{k}$, and then optimizes the observation set $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ based on linear programming. We will demonstrate that the adversary cannot reduce its uncertainty of the public state by reversing the proposed filtering policy.
Figure 2: The structure of privacy filter.
4.2 Randomization
We consider the following random set
${\mathcal{S}}_{k|k}^{x}=\left[\begin{array}[]{c}x_{k}-{\alpha}_{k}\left(x_{k}-\underline{X}_{k|k-1}\right)\\ x_{k}+{\beta}_{k}\left(\overline{X}_{k|k-1}-x_{k}\right)\\ \end{array}\right],$ (41)
where $\alpha_{k}$ and $\beta_{k}$ are uniform random variables with
${\alpha}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right],$
and
${\beta}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right].$
Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius between the actual public state $X_{k}=x_{k}$ and the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius between $x_{k}$ and the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ is a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state. Also, we have $\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ as
$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)$
$=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
$\leqslant$ $\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}$
$+\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
$=$ $\epsilon^{x}.$
We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to enhance the privacy level.
4.3 Privacy Filter Optimization
[…]
$\mathbf{P_{1}}:$ $\max_{\mathcal{M}^{x}_{k|k}}\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ (42)
[…]
where $\mathcal{S}^{x}_{k|k}\subseteq\mathcal{M}^{x}_{k|k}$ is required by the structure of the privacy filter, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ describes the prior space of $\mathcal{M}^{x}_{k|k}$, and $\mathrm{Vol}\left({\mathcal{M}}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ enforces the requirement of data utility. We next demonstrate that the optimization problem $\mathbf{P_{1}}$ can be solved via linear programming.
Theorem 7.
The privacy filter optimization problem $\mathbf{P_{1}}$ is equivalent to the following linear programming
$\mathbf{P_{2}}:\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$ (48)
[…]
$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right.,\text{for}\;Z=X,Y$ (59)
where $\epsilon^{y}\geqslant 0$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$ and $p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}\in\mathcal{R}^{n_{x}}$.
Proof 4.1.
See Appendix D
Figure 3: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.01$.
Figure 4: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.5$.
As a result, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3. Although the linear programming problem $\mathbf{P_{2}}$ is deterministic, we can demonstrate that the adversary cannot reduce its uncertainty by reversing the optimization process.
Proposition 8.
If the adversary infers the private state via the inference attack (2)-(32), then it cannot find a smaller interval subset of $\mathcal{M}_{k|k}^{x,\star}$ containing the actual public state $X_{k}=x_{k}$ via reverting the linear programming.
Proof 4.2.
On the one hand, the actual public state $X_{k}=x_{k}$ can be any element of $\mathcal{X}_{k|k-1}$, and $S_{k}^{x}$ is a random subset of $\mathcal{X}_{k|k-1}$ containing $x_{k}$. Thus, the upper or lower endpoint of $S_{k}^{x}$ may coincide with $x_{k}$.
On the other hand, since $\mathcal{M}_{k|k}^{x,\star}$ is the optimal observation set obtained from the linear programming, for any $S_{k}^{x}$ in the vicinity of $\mathcal{M}_{k|k}^{x,\star}$, i.e., $S_{k}^{x}\in\left\{S_{k}^{x}\left|\lim_{\left\|\Delta\right\|_{1}\rightarrow 0}{S_{k}^{x}+\Delta=\mathcal{M}_{k|k}^{x,\star}}\right.\right\}$, the optimal solution remains $\mathcal{M}_{k|k}^{x,\star}$.
As a result, it is possible that the upper or lower endpoint of $\mathcal{M}_{k|k}^{x,\star}$ corresponds to the actual public state. Thus, the minimal interval observation set for the adversary is $\mathcal{M}_{k|k}^{x,\star}$.
According to Proposition 8, even if the adversary knows the structure of the privacy filter, it cannot revert the linear programming to obtain a smaller interval containing $x_{k}$ to reduce uncertainty. In other words, the privacy filer is robust against the reverse optimization attack.
5 Numerical Verification
In this section, we study the performance of privacy filter for the production-inventory problem with the following parameters
$A_{1}=\left[\begin{matrix}1.00&0.00\\ 0.00&1.00\\ \end{matrix}\right],A_{2}=\left[\begin{matrix}0.40&0.80\\ 0.60&0.20\\ \end{matrix}\right],$
$A_{3}=\left[\begin{matrix}0.50&-0.90\\ -0.10&-0.10\\ \end{matrix}\right],A_{4}=\left[\begin{matrix}-0.10&-0.90\\ 0.10&0.00\\ \end{matrix}\right],$
$B_{1}=\left[\begin{matrix}-1.00&0.00\\ 0.00&-1.00\\ \end{matrix}\right],B_{2}=\left[\begin{matrix}4.20&0.00\\ 0.00&2.40\\ \end{matrix}\right],$
$\left(\mathcal{W}_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\right],$
$\left(\mathcal{W}_{k}^{y}\right)^{\top}=\left[\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\right].$
The initial state sets are assumed to be
$\left(\mathcal{X}_{0|-1}\right)^{\top}=\left[\begin{matrix}1.00&0.24&1.20&0.40\\ \end{matrix}\right],$ (60)
$\left(\mathcal{Y}_{0|-1}\right)^{\top}=\left[\begin{matrix}2.40&0.60&3.70&1.30\\ \end{matrix}\right].$ (61)
In our simulation, the initial states are uniformly sampled from the bounded sets (60)-(61). To simulate the approximate periodic fluctuations in demand and productivity, the actual process noises are set to be
$\left(W_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.88+0.03\cos\left(\frac{2\pi k}{30+7\rho_{k}}\right)&1.94\\ \end{matrix}\right],$
$\!\!\left(W_{k}^{y}\right)\!^{\top}\!=\!\left[\!\begin{matrix}0.944\!+\!0.006\cos\left(\!\frac{2\pi k}{7+2\gamma_{k}}\!\right)\!\!&\!\!0.33+0.094\sin\left(\!\frac{2\pi k}{7+4\tau_{k}}\!\right)\\ \end{matrix}\!\right]\!,$
where $\rho_{k}$, $\gamma_{k}$ and $\tau_{k}$ are uniform random variables in $\left[0,1\right]$. As discussed in Section 2, the production rate is private but the inventory information has to be released.
We first plot the trajectories of system states and their interval tubes in Fig.3 and Fig.4 under the optimal privacy filter design for different values of $\epsilon^{x}$. We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$. The pink areas in these figures represent the uncertainty sets of system states. As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate. However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s estimation of the production rate to become less accurate. Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
Figure 5: The truncated Gaussian mechanism.
We analyze the utility-privacy trade-off performance of the optimal privacy filtering policy and compare it with the truncated Gaussian mechanism, which is proposed in [18] for differential privacy. In the truncated Gaussian mechanism, illustrated in Fig. 5, the original state $x_{k}$ is perturbed by additive noise $v_{k}$ drawn from a truncated Gaussian distribution defined over the interval $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, with zero mean and variance $\left(\epsilon^{x}\right)^{2}$. Subsequently, the perturbed observation set is released as follows:
$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ \end{array}\right].$ (62)
Since the noise $v_{k}$ is constrained within $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, the released public state $x_{k}$ is guaranteed to reside within the interval $\mathcal{M}_{k|k}^{x}$, and the utility of $x_{k}$ is quantified by $\epsilon^{x}$.
The privacy-utility trade-off is evaluated by plotting the privacy level of the production rate against the utility of the inventory under different approaches, as depicted in Fig. 6. The results indicate that increasing the utility of inventory leads to a reduction in the privacy level of the production rate, highlighting the inherent trade-off between data utility and privacy protection. Compared to the truncated Gaussian mechanism, the optimal privacy filter achieves a higher privacy level while maintaining lower data distortion.
Figure 6: The privacy level of the private state and utility of the public state.
6 Conclusion
In this paper, we developed a volumetric privacy measure for dynamic systems with UBB noise. We defined the inference attack that an adversary uses to estimate the private state and introduced a volumetric measure to evaluate the privacy level. We provided computational approaches based on interval analysis and analyzed the theoretical properties of the proposed measure. Furthermore, we proposed an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method was demonstrated through a production-inventory case study.
Appendix A Proof of Lemma 2
At the time step $k=0$, the adversary only has prior knowledge, i.e., $\mathcal{Y}_{0|-1}$, therefore, its inference set is $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$.
To prove Lemma 2 for $k\geqslant 1$, we need the following lemma that computes the tightest interval by forward reachability analysis.
Lemma 9.
[1, 2] Consider the static system $S=AM+BW$, where $M$ and $W$ are bounded intervals, the tightest interval for $S$, i.e., $\mathcal{S}$ can be computed as
$\mathcal{S}=\Psi\left(A\right)\mathcal{M}+\Psi\left(B\right)\mathcal{W},$ (63)
where
$\Psi\left(N\right)=\left[\begin{matrix}\frac{N+\left|N\right|}{2}&\frac{N-\left|N\right|}{2}\\ \frac{N-\left|N\right|}{2}&\frac{N+\left|N\right|}{2}\\ \end{matrix}\right].$ (64)
Also, we can compute the radius and center of the intervl $\mathcal{S}$ via
$p^{s}=\left|A\right|p^{m}+\left|B\right|p^{w},$ (65)
$c^{s}=Ac^{m}+Bc^{w}.$ (66)
When $A_{1}$ and $A_{2}$ are invertible, we have
$X_{k-1}=A_{1}^{-1}X_{k}+\left(-A_{1}^{-1}A_{2}\right)Y_{k-1}+\left(-A_{1}^{-1}B_{1}\right)W_{k}^{x},$
and
$Y_{k-1}=A_{2}^{-1}X_{k}+\left(-A_{2}^{-1}A_{1}\right)X_{k-1}+\left(-A_{2}^{-1}B_{1}\right)W_{k}^{x}.$
According to Lemma 9, the tightest intervals for (2.3) and (2.3) are (2) and (2). Then, the intersection of different intervals, i.e., (12) and (13), can be computed with (24) and (27). Finally, the one-step forward reachable set (14) can be approximated with the tightest interval (32) based on Lemma 9, and the calibrated uncertainty set $\mathcal{X}_{k|k}$ and the tightest prior inference set $\mathcal{Y}_{k|k-1}$ can be approximated similarly. Finally, since at the time step $k=0$, the backward calibration (2.3) and (2.3) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (36).
Appendix B Proof of Lemma 3
According to Lemma 9, the radius of $\mathcal{M}^{y}_{k-1|k}$ can be computed as
$p_{k-1|k}^{m,y}=\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}+\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x},$
where $p^{m,x}_{k|k}$, $p_{k-1|k-1}^{x}$ and $p^{w,x}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{X}_{k-1|k-1}$ and $\mathcal{W}_{k}^{x}$, respectively. Since $\mathcal{Y}_{k-1|k}$ is the intersection result from $\mathcal{M}^{y}_{k-1|k}$ and $\mathcal{Y}_{k-1|k-1}$, the radius of $\mathcal{Y}_{k-1|k}$ is smaller than the radius of $\mathcal{M}^{y}_{k-1|k}$, i.e., $p_{k-1|k}^{y}\leqslant p_{k-1|k}^{m,y}$. Also, the radius of $\mathcal{Y}_{k|k}$ can be computed as
$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$
Therefore, with (B), we have
$p_{k|k}^{y}\leqslant\left|A_{3}\right|p_{k-1|k}^{x}+\left|B_{2}\right|p_{k}^{w,y}+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x}$
$+\left|A_{4}\right|\left(\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}\right).$
Since $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$ and $\mathcal{X}_{k-1|k}$ is a subset of $\mathcal{M}_{k-1|k-1}^{x}$, we have $p_{k-1|k}^{x}\leqslant p^{m,x}_{k-1|k-1}\leqslant\overline{p}^{x}$ for any $k\geqslant 1$, thus we have (3).
Appendix C Proof of Theorem 5
The difference set $\Delta\mathcal{Y}_{k|k}$ is computed as,
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}+\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
where
$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}-\mathcal{M}_{k-1|k}^{x}$
$=\left[\begin{array}[]{c}\min\left\{\underline{X}_{k-1|k-1}-\underline{M}_{k-1|k}^{x},0\right\}\\ \max\left\{\overline{X}_{k-1|k-1}-\overline{M}_{k-1|k}^{x},0\right\}\\ \end{array}\right],$ (69)
$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$
$=\left[\begin{array}[]{c}\min\left\{\underline{Y}_{k-1|k-1}-\underline{M}_{k-1|k}^{y},0\right\}\\ \max\left\{\overline{Y}_{k-1|k-1}-\overline{M}_{k-1|k}^{y},0\right\}\\ \end{array}\right].$ (72)
Therefore, the volume of the difference set is (38).
With Lemma 9, we have
$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$ (73)
where the radius $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ can be computed via
$2p_{k-1|k}^{\Delta z}$
$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}-\min\left\{\underline{Z}_{k-1|k-1}-\underline{M}_{k-1|k}^{z},0\right\}$
$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}+\max\left\{\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},0\right\}$
$=$ $\max\left\{0,\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}+\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},\right.$
$\left.\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\right\},\text{for}\,Z=X,Y,$ (74)
which satisfies $p_{k-1|k}^{\Delta z}\geqslant 0$. As a result, we have
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
$=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
=$\overset{(a)}{=}\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
$\leqslant\left\|A_{3}\right\|\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
$=\left\|A_{3}\right\|\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$ (75)
where $(a)$ is due to $p_{k-1|k}^{\Delta x}\geqslant 0$ and $p_{k-1|k}^{\Delta y}\geqslant 0$.
Besides, given an interval $\mathcal{X}$, we can express it with its center point and radius, i.e., $\mathcal{X}=\left[\begin{array}[]{c}\frac{c-p}{2}\\ \frac{c+p}{2}\end{array}\right]$. Therefore, we have (C),
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
$=$ $\left\|\overline{Y}_{k|k-1}-\overline{Y}_{k|k}\right\|_{1}+\left\|\underline{Y}_{k|k-1}-\underline{Y}_{k|k}\right\|_{1}$
$=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
$\geqslant$ $\left\|\overline{Y}_{k|k}+\underline{Y}_{k|k}-\left(\overline{Y}_{k|k-1}+\underline{Y}_{k|k-1}\right)\right\|_{1}$
$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$ (76)
Appendix D Proof of Theorem 7
To maximize the privacy level, it is equivalent to minimize the amount of uncertainty reduction since we have $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, where the prior uncertainty set $\mathcal{Y}_{k|k-1}$ is fixed at time step $k$.
Besides, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=\left\|p_{k|k}^{\Delta y}\right\|_{1}=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$, where the elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ are non-negative vectors as shown in (C). Therefore, we can replace the objective function with the slack variable $\epsilon^{y}$ and add $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\leqslant\epsilon^{y}$ as a new constraint, and then minimize $\epsilon^{y}$.
Since $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (24) and (27) with the constraints of difference sets (C). Also, the objective function increases with any elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ since the elements of $p_{k-1|k}^{\Delta x}$, $p_{k-1|k}^{\Delta y}$, $\left|A_{3}\right|$ and $\left|A_{4}\right|$ are non-negative. As a result, we can replace the constraint of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, i.e., (C), with inequalities (59), and let $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ be decision variables.
Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.
Finally, the objective and the constraints are linear functions of the decision variables, thus, the optimal privacy filter can be obtained by solving the linear programming $\mathbf{P_{2}}$.