### v4

# A Volumetric Privacy Measure for Dynamical Systems With Bounded Disturbance

Abstract
This paper presents a volumetric privacy framework for dynamical systems subject to bounded disturbances, developed without requiring prior knowledge of their probability distributions. We consider systems with both public and private states, where a set containing the public state is shared as the observation. An adversary is assumed to execute an inference attack by exploiting the observed public state set to estimate an uncertainty set for the private state. The volume of this inferred set quantifies the adversary’s estimation uncertainty and serves as the proposed volumetric privacy metric. Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that it is bounded by the information gain from the observation set. Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the volumetric privacy level. The effectiveness of the proposed approach is validated through a production–inventory case study. Results show that the optimal privacy filter significantly improves robustness against inference attacks and outperforms two baseline mechanisms based on additive noise and quantization.
keywords:
Volumetric privacy measure; privacy protection; interval analysis; bounded disturbance.
thanks: The work was partially supported by the Research Grants Council of Hong Kong under Project CityU 21208921, a grant from Chow Sang Sang Group Research Fund sponsored by Chow Sang Sang Holdings International Limited.
1 Introduction
1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information. Specifically, correlations between shared metrics and underlying operational information can be exploited by adversaries to develop competitive and malicious strageties. This challenge highlights the critical need for methodologies that preserve data utility while ensuring privacy protection for dynamic systems.
Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information, e.g., inventory–production systems and connected vehicles. However, the notion of privacy in such systems remains insufficiently explored. Within the context of set-membership estimation, the states of these systems can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty.
Motivated by these considerations, this paper investigates the notion of volumetric privacy for systems affected by bounded disturbance. We develop privacy-preserving strategies aimed at maximizing an adversary’s uncertainty, i.e., the volume of uncertainty set, in inferring private states. The proposed approaches are applicable to both deterministic and stochastic systems, without requiring prior knowledge of the underlying probability distributions.
1.2 Related Work
Stochastic approaches to privacy primarily include differential privacy and information-theoretic methods. Differential privacy (DP) [1] has been incorporated into dynamic settings through differentially private Kalman filtering [2], DP-preserving average consensus via noise injection [3], and minimal-noise mechanisms for multi-agent systems based on observability properties [4]. Recent work [5] introduced a trace-based variance–expectation ratio to quantify topology preservation and derived optimal noise designs, while [6] provided a comprehensive survey of DP in dynamical systems. In parallel, information-theoretic approaches quantify privacy leakage using mutual information or conditional entropy. Recent studies include mutual-information-based private filtering for hidden Markov models [7], directed-information-based privacy filters for linear systems [8, 9], and recent extensions to partially observable Markov decision processes addressing privacy-aware estimation and control [10, 11].
Most existing studies focus on deterministic or stochastic systems with unbounded noise and known distributions, leaving privacy protection for systems subject to unknown-but-bounded disturbances relatively unexplored. Recent works [12, 13] developed differentially private set-based estimators using truncated noise, while [14] introduced guaranteed privacy concepts and optimization methods for $\mathcal{H}_{\infty}$-based privacy-preserving interval observers. State-opacity-based methods [15, 16] ensured indistinguishable outputs between secret and non-secret states but did not quantify the associated estimation uncertainty. Note that set-membership estimators typically characterize uncertainty through bounded geometric sets such as intervals [17], zonotopes [18], or ellipsoids [19], where the corresponding set volume naturally describes the amount of estimation uncertainty. Motivated by this, we analyze privacy leakage in systems with private and public states and propose a volumetric approach that maximizes the private-state set volume, thereby enhancing privacy while explicitly accounting for geometric effects under inference attacks.
There are some related deterministic approaches to privacy without adding noise, e.g., plausible deniability [20] and noiseless privacy [21]. In [20], privacy leakage in deterministic systems was measured by the volume of reachable state sets, and was determined by the observability. However, this framework does not apply to systems with bounded disturbance, where inference uncertainty depends on both observability and disturbance. Moreover, the problem of privacy filter design was not addressed in [20], whereas we propose a concrete design using randomization and optimization. The work in [22] addressed parameter privacy in deterministic systems via constrained convex generators (CCGs), which differs from our private state protection setting. Also, defense strategies in [22] involve ceasing information sharing or altering parameters, which might be unsuitable for fixed-parameter systems with continuous communication. As discussed in Sec. 3.1, the complexity of CCG-based inference grows exponentially with time, motivating our use of interval analysis for computational efficiency.
Noiseless privacy [21] and non-stochastic privacy [23] employed non-stochastic information-theoretic approaches to limit information leakage, assuming static private-variable domains and without accounting for temporal dependencies in sequential data. While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs. However, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation. Building on this insight, the proposed volumetric privacy filter dynamically evaluates the private state set and adapts the observation accordingly, thereby achieving higher privacy levels with lower data distortion, as shown in Sec. 5.
Finally, other deterministic privacy-preserving approaches often exploit observability reduction or state decomposition to protect private information in multi-agent systems. For example, the authors in [24] established a connection between network privacy and its observability space, proposing a privacy-aware communication protocol that achieves average consensus while protecting initial states. In [3], the trace of the observability Gramian was employed to quantify information leakage through an intruder node, and an online optimization strategy was proposed to adapt communications in order to minimize such leakage. Privacy-preserving consensus designs were also proposed in [25, 26], including augmented states and novel consensus algorithms with simultaneous accuracy, resilience, and privacy guarantees. State decomposition methods, as in [27], split each node’s state into randomized components to prevent disclosure of individual states during consensus.
1.3 Contributions
This paper studies privacy for dynamic systems with bounded disturbance as shown in Fig. 1, where the system state is split into public $X_{k}$ and private $Y_{k}$, and an adversary uses the public observation set $\mathcal{M}_{k|k}^{x}$ to infer $Y_{k}$ via an uncertainty set $\mathcal{Y}_{k|k}$.
Figure 1: The system setup.
The primary contribution of this work is the development of an extensible framework for privacy analysis and mitigation in dynamic systems subject to bounded disturbance. This contribution can be summarized in three principal aspects. (1) Volumetric Privacy Measure: We introduce a privacy metric based on the volume of the estimated private-state set obtained via set-membership estimation given the available observations. (2) Privacy Level Computation: We develop computational methods to quantify privacy level and prove that the relevant privacy leakage is bounded by the information gain from the observations. (3) Optimal Privacy Filter: Since the inappropriate choice of the observation set would lead to privacy leakage of the private state, we design a randomized, optimization‐based filter that perturbs and then refines observations to maximize inference uncertainty of attackers. Finally, the proposed framework is demonstrated on a production–inventory case study, showing that our privacy filter significantly reduces the adversary’s capability to estimate the private production rate.
1.4 Outline
The rest of the paper is organized as follows. Section 2 introduces the system model, the inference attack and defines the volumetric privacy and utility measure. Section 3 provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure. Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level. Section 5 presents numerical results, followed by the conclusions in Section 6.
1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. For the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z\!+\!R|Z\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$ with the absolute value $\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The 1-norm of the matrix $A$ is defined as $\left\|A\right\|_{1}\!\frac{\Delta}{=}\!\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\!\times\!n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$. The operator $diag(v)$ denotes a diagonal matrix constructed from the vector $v$.
2 System Model and Inference Attack
2.1 System Model
We consider the following stable system model $\mathbf{G_{1}}$,
$\mathbf{G_{1}}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ \end{array}\right$., (3)
where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the unknown disturbance with bounded sets $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$. Note that the underlying probability distributions of $W_{k}^{x}$ and $W_{k}^{y}$ are unknown. Besides, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively. We assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
2.2 Motivating Examples
In this subsection, we consider two motivating examples to illustrate the necessity of protecting privacy of systems $\mathbf{G_{1}}$.
Production-inventory system: In supply chain management [28, 29], the inventory level $X_{k}$ and the production rate $Y_{k}$ evolves according to $\mathbf{G}_{1}$. While firms may disclose inventory information $X_{k}$ to distributors to boost sales, the production rate $Y_{k}$ contains sensitive strategic information such as production efficiency and supply chain operations. Since $X_{k}$ and $Y_{k}$ are correlated, releasing $X_{k}$ directly risks revealing private production details. Therefore, it is necessary to transform or mask observations to preserve the privacy of $Y_{k}$ while maintaining the utility of public inventory data $X_{k}$.
Traffic management system: In intelligent transportation, vehicles may report their velocities to a central controller to optimize traffic flow, e.g., by adjusting the speed limit on highways. Given bounded disturbances from environmental factors like uneven ground, the vehicle dynamics fit the model $\mathbf{G}_{1}$ with unknown-but-bounded disturbance. Here, velocity $X_{k}$ can be considered public data used for traffic management, while position $Y_{k}$ is private, as it can be used to identify individual vehicles. To protect location privacy, vehicles may intentionally report blurred or randomized velocity observations that preserve system utility but reduce the risk of precise location inference.
2.3 Inference Attack
Due to the presence of unknown disturbance terms belonging to bounded sets, multiple private states may correspond to the same public state. Consequently, for a given public state set, an adversary cannot determine the exact private state; instead, it can only identify a corresponding uncertainty set. we next define how the adversary infers the private state using set-theoretic operations.
We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$ and other elements to obfuscate the adversary’s estimation of private states. Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set. This process is referred to as the inference attack. We next define the inference attack recursively.
At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of states can be predicted based on the system model (3), i.e.,
$\mathcal{X}_{k|k-1}$ $=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x},$ (4)
$\mathcal{Y}_{k|k-1}$ $=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (5)
After receiving the observation set of the public state $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$, the adversary extracts new information from $\mathcal{M}^{x}_{k|k}$ and updates the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via the following steps,
$\!\!\!\!\mathcal{M}_{k\!-\!1|k}^{x}\!\!=\!A_{1}^{\!-\!1}\!\mathcal{M}_{k|k}^{x}$ $\!\oplus\!\left(\!-\!A_{1}^{\!-\!1}\!A_{2}\!\right)\!\mathcal{Y}_{k\!-\!1|k\!-\!1}\!\oplus\!\left(-\!A_{1}^{\!-\!1}\!B_{1}\!\right)\!\mathcal{W}_{k}^{x}\!,\!\!$ (6)
$\!\!\!\!\mathcal{M}_{k\!-\!1|k}^{y}\!=\!A_{2}^{\!-\!1}\!\mathcal{M}_{k|k}^{x}$ $\!\oplus\!\left(\!-A_{2}^{\!-\!1}\!A_{1}\!\right)\!\mathcal{X}_{k\!-\!1|k\!-\!1}\!\oplus\!\left(\!-A_{2}^{\!-\!1}\!B_{1}\!\right)\!\mathcal{W}_{k}^{x}\!,\!\!$ (7)
$\mathcal{X}_{k-1|k}$ $=\mathcal{M}_{k-1|k}^{x}\cap\mathcal{X}_{k-1|k-1},$ (8)
$\mathcal{Y}_{k-1|k}$ $=\mathcal{M}_{k-1|k}^{y}\cap\mathcal{Y}_{k-1|k-1},$ (9)
where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (6) and (7), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (8) and (9).
According to the system dynamics (3), the adversary estimates the uncertainty set of $Y_{k}$ via the following forward inference,
$\mathcal{Y}_{k|k}=A_{3}\mathcal{X}_{k-1|k}\oplus A_{4}\mathcal{Y}_{k-1|k}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (10)
Finally, the public state set can be further calibrated via,
$\mathcal{X}_{k|k}=\mathcal{M}^{x}_{k|k}\cap\mathcal{M}^{x}_{k|k-1},$ (11)
$\mathcal{M}^{x}_{k|k-1}=$ $A_{1}\mathcal{X}_{k-1|k}\oplus A_{2}\mathcal{Y}_{k-1|k}\oplus B_{1}\mathcal{W}_{k}^{x},$ (12)
where $\mathcal{M}^{x}_{k|k-1}$ is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (6)-(9), and the forward inference (10)-(12). The backward calibration (6)-(9) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
Proposition 1.
For any $k\geqslant 1$, the adversary’s uncertainty set for the private state (10) is a subset of its corresponding prediction set (5), i.e., $\mathcal{Y}_{k|k}\subseteq\mathcal{Y}_{k|k-1}$. Moreover, given uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ that contain the true system states $X_{k-1}\!=\!x_{k-1}$ and $Y_{k-1}\!=\!y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the true public state $X_{k}\!=\!x_{k}$, then the inference set $\mathcal{Y}_{k|k}$ contains the true private state $Y_{k}\!=\!y_{k}$.
Proof 2.1.
Since $\mathcal{M}^{x}_{k|k}$ contains $x_{k}$, it follows from (6) that $x_{k-1}\!\in\!\mathcal{M}^{x}_{k-1|k}$. As $x_{k-1}$ also belongs to $\mathcal{X}_{k-1|k-1}$, their intersection $\mathcal{X}_{k-1|k}$ necessarily contains $x_{k-1}$. By the same reasoning, $y_{k-1}\!\in\!\mathcal{Y}_{k-1|k}$. Propagating through the system dynamics $\mathbf{G}_{1}$ yields $y_{k}\!\in\!\mathcal{Y}_{k|k}$. Finally, since $\mathcal{X}_{k-1|k}\!\subseteq\!\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k}\!\subseteq\!\mathcal{Y}_{k-1|k-1}$, we have $\mathcal{Y}_{k|k}\!\subseteq\!\mathcal{Y}_{k|k-1}$.
According to Proposition 1, the adversary can reduce its uncertainty of the private state via the inference attack since it can obtain a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$. In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain the actual private state.
2.4 Privacy and Utility Measures
As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$. The attacker has more inference uncertainty about the private state if $\mathcal{Y}_{k|k}$ contains more elements. However, since the state space is continuous, the number of elements in such sets is uncountable. To address this, we propose to use the volume of the uncertainty set as a quantitative measure of privacy. Specifically, we define
$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\int_{\mathcal{Y}_{k|k}}\mathrm{d}y,$ (13)
where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure (i.e., the geometric volume) of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n}$. We then define the privacy measure at time $k$ as
$P_{k}\left(\mathcal{Y}_{k|k}\right):=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right).$ (14)
Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero. To protect the system from inference attack, we can increase the privacy level of the private state via maximizing $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$.
On the other hand, it causes more utility distortion of the public state if the observation set $\mathcal{M}^{x}_{k|k}$ contains more elements, since it is more difficult for the receiver to recover the actual public state. To address this, we would reduce distortion via maximizing the following defined utility of public state set $\mathcal{M}^{x}_{k|k}$,
$U_{k}\left(\mathcal{M}^{x}_{k|k}\right)=1/\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right).$ (15)
Notably, set operations in the proposed inference attack and the volume computation are generally computational costly due to the continuous nature of the state space. Consequently, it is essential to provide efficient computational tools for privacy level evaluation. Also, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage. To address these challenges, we study computation approaches for inference attack in Sec. 3, and present an optimal privacy filter design to reduce privacy leakage in Sec. 4.
3 Inference Attack Approximation
As addressed in existing set-membership estimation approaches [18, 19], the set operations involved in inference attacks can be computationally expensive. To mitigate this complexity, uncertainty sets are often restricted to specific geometric forms, enabling more efficient implementation of set operations. However, more complex representations generally incur higher computational costs in volume evaluation. In the following, we present two approximation methods for implementing inference attacks and analyze their computational complexity, based on which we establish properties of volumetric privacy.
3.1 Inference Attack Approximation via CCGs
In this subsection, we show that the inference attack can be approximated using the constrained convex generator (CCG), a general set representation proposed in [30].
Definition 2.
[30] The constrained convex generator $\mathcal{Z}=\left(G,c,A,b,\mathcal{C}\right)\subset\mathcal{R}^{n}$ is defined as
$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$ (16)
where $G\in\mathcal{R}^{n\times n_{g}}$, $c\in\mathcal{R}^{n}$, $A\in\mathcal{R}^{n_{c}\times n_{g}}$, $b\in\mathcal{R}^{n_{c}}$ and $\mathcal{C}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{n_{p}}\right\}$, and $\mathcal{C}_{i}\subset\mathcal{R}^{m_{i}}$ are convex sets with $\sum_{i=1}^{n_{p}}{m_{i}}=n_{g}$.
The CCG encompasses a wide range of useful set representations, including zonotopes, ellipsoids, and intervals [30]. Moreover, common set operations such as the Minkowski sum and intersection admit analytical expressions, enabling its application to approximate the inference attack.
Proposition 3.
[30] Given CCGs $\mathcal{X}\!=\!\left(G_{x},c_{x},A_{x},b_{x},\mathcal{C}_{x}\right)\!\subset\!\mathcal{R}^{n}$ and $\mathcal{Y}\!=\!\left(G_{y},c_{y},A_{y},b_{y},\mathcal{C}_{y}\right)\!\subset\!\mathcal{R}^{n}$, and a matrix $R\in\mathcal{R}^{m\times n}$, we have
$R\mathcal{X}=\left(RG_{x},Rc_{x},A_{x},b_{x},\mathcal{C}_{x}\right),$
$\mathcal{X}\!\oplus\!\mathcal{Y}$ $\!=\!\left(\!\left[\begin{matrix}G_{x}&\!G_{y}\\ \end{matrix}\right]\!,c_{x}\!+\!c_{y},\mathrm{diag}\left(\!\left[\begin{matrix}A_{x}&\!A_{y}\\ \end{matrix}\right]\!\right)\!,\left[\!\begin{array}[]{c}b_{x}\\ b_{y}\\ \end{array}\!\right]\!,\left\{\mathcal{C}_{x},\mathcal{C}_{y}\right\}\!\!\right)\!,$ (19)
$\mathcal{X}\cap\mathcal{Y}$ $=\left(\left[\begin{matrix}G_{x}&0\\ \end{matrix}\right],c_{x},\left[\begin{matrix}A_{x}&0\\ 0&A_{y}\\ G_{x}&-G_{y}\\ \end{matrix}\right],\left[\begin{array}[]{c}b_{x}\\ b_{y}\\ c_{y}-c_{x}\\ \end{array}\right],\left\{\mathcal{C}_{x},\mathcal{C}_{y}\right\}\right).$ (23)
According to the computation rules in Proposition 3, the inference attack from (6) to (12) can be directly implemented, if we assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as CCGs. However, as shown in the next lemma, the computational complexity of CCG-based inference grows exponentially over time.
Proposition 4.
The computational complexity of the CCG-based inference attack at time $k$ is at least $\mathcal{O}\!\left(c^{k-1}n^{3}\right)$ for some constant $c>1$, and both the column dimension of the generator matrices $G$ and the number of constraints grow exponentially with $k$.
Proof 3.1.
The dominant operation in the inference attack from (6) to (12) is the multiplication of an $n\times n$ matrix with an $n\times m$ matrix, which has computational complexity $\mathcal{O}(mn^{2})$. For simplicity, we assume that at time $k$ the sets $\mathcal{X}_{k-1|k-1}$, $\mathcal{Y}_{k-1|k-1}$, $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, and $\mathcal{M}^{x}_{k|k}$ all have generator matrices $G$ of size $n\times n$ and are described by $n$ constraints.
According to Proposition 3, after one inference step, the Minkowski sum and intersection operations cause the generator matrix in $\mathcal{X}_{k|k}$ to grow to size $n\times(c\,n)$ for some constant $c>1$, while the number of constraints increases by a factor $d>1$. Thus, both the column dimension of $G$ and the number of constraints grow exponentially with $k$. Consequently, due to the exponentially increasing column dimension, the computational complexity of matrix multiplications in the inference attack at time $k$ is at least $\mathcal{O}(c^{k-1}n^{3})$.
The high computational complexity of CCG-based inference renders real-time implementation of the inference attack and the privacy filter design in Sec. 4 intractable for large $k$. Order-reduction techniques can be employed to reduce this complexity, albeit at the cost of some loss in inference accuracy. However, such techniques also complicate the analysis of the proposed volumetric privacy metric. For clarity and focus, we defer a detailed discussion of these techniques to future work.
3.2 Inference Attack Approximation via Interval Analysis
We next consider an interval-based approximation approach for computing the inference sets. This approach is a special case of CCG-based inference but significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic. Moreover, interval-based inference serves as the foundation for designing the optimal privacy filter, as discussed in Sec. 4.
Definition 5.
The interval $\mathcal{X}=\{X\mid\underline{X}\leqslant X\leqslant\overline{X}\}$ is defined as $\begin{bmatrix}\underline{X}\\[3.00003pt] \overline{X}\end{bmatrix}$ with the lower and upper bounds $\underline{X}$ and $\overline{X}$. An interval $\mathcal{X}$ can also be expressed as a special case of the CCG representation $\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathcal{R}^{n_{x}},\;\|\xi\|_{\infty}\leqslant 1\right\}$, with the center point $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and the radius $p^{x}=\frac{\overline{X}-\underline{X}}{2}$. Also, the volume of the interval $\mathcal{X}$ is computed as $\mathrm{Vol}\left(\mathcal{X}\right)=\sum_{i=1}^{n}{\left(\overline{X}\left(i\right)-\underline{X}\left(i\right)\right)}$ where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension of the interval $\mathcal{X}$, respectively.
Given a block matrix $A=[A_{1},\,A_{2}]$, the multiplication of $A$ with an interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$. If $\mathcal{X}$ and $\mathcal{Y}$ are intervals, their Minkowski sum is $\mathcal{X}\oplus\mathcal{Y}=\begin{bmatrix}\underline{X}+\underline{Y}\\[3.00003pt] \overline{X}+\overline{Y}\end{bmatrix}$, and their difference, used only for volume evaluation, is $\mathcal{X}\setminus\mathcal{Y}=\begin{bmatrix}\underline{X}-\underline{Y}\\[3.00003pt] \overline{X}-\overline{Y}\end{bmatrix}$. The intersection of $\mathcal{X}$ and $\mathcal{Y}$ is $\mathcal{X}\cap\mathcal{Y}=\begin{bmatrix}\max\{\underline{X},\underline{Y}\}\\[3.00003pt] \min\{\overline{X},\overline{Y}\}\end{bmatrix}$.
We now assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as intervals. Under this assumption, the interval-based inference attack can be implemented using the following lemma.
Lemma 6.
The recursive inference interval from (6) to (9) can be computed via
$\mathcal{M}_{k-1|k}^{x}$ $=\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}\oplus\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
$\oplus\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (24)
$\mathcal{M}_{k-1|k}^{y}$ $=\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}\oplus\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{X}_{k-1|k-1}$
$\oplus\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (25)
$\mathcal{X}_{k-1|k}$ $=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{x},\underline{X}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{x},\overline{X}_{k-1|k-1}\right\}\\ \end{array}\right],$ (28)
$\mathcal{Y}_{k-1|k}$ $=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{y},\underline{Y}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{y},\overline{Y}_{k-1|k-1}\right\}\\ \end{array}\right],$ (31)
$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!$ $\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!\oplus\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!\oplus\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (32)
$\mathcal{X}_{k|k}$ $=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k|k}^{x},\underline{M}_{k|k-1}^{x}\right\}\\ \min\left\{\overline{M}_{k|k}^{x},\overline{M}_{k|k-1}^{x}\right\}\\ \end{array}\right],$ (35)
$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}\!=\!\Psi\!$ $\left(A_{3}\right)\!\mathcal{X}_{k-1|k}\!\oplus\!\Psi\!\left(A_{4}\right)\mathcal{Y}_{k-1|k}\!\oplus\!\Psi\!\left(B_{2}\right)\mathcal{W}_{k}^{y}\!,\!$ (36)
with
$\Psi\left(\star\right)=\left[\begin{matrix}\frac{\star+\left|\star\right|}{2}&\frac{\star-\left|\star\right|}{2}\\ \frac{\star-\left|\star\right|}{2}&\frac{\star+\left|\star\right|}{2}\\ \end{matrix}\right].$
Also, the prior inference set of $Y_{k}$ is
$\!\!\!\!\!\mathcal{Y}_{k|k\!-\!1}\!=\!\Psi\!\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!\oplus\!\Psi\!\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!\oplus\!\Psi\!\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y}\!,\!\!$ (37)
if $k\geqslant 1$. If $k=0$, then $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$ and
$\mathcal{X}_{0|0}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{0|0}^{x},\underline{X}_{0|-1}\right\}\\ \min\left\{\overline{M}_{0|0}^{x},\overline{X}_{0|-1}\right\}\\ \end{array}\right].$ (40)
Proof 3.2.
See Appendix A.
Given the interval-based inference approach described in Lemma 6, the computational complexity of the inference attack can be characterized as follows.
Proposition 7.
The computational complexity of the inference attack via interval analysis is $\mathcal{O}(n^{3})$.
Proof 3.3.
The dominant operation in the inference attack involves matrix multiplication. Since the matrix $\Psi(\star)$ has dimensions $(2n\times 2n)$, the corresponding computational complexity is $\mathcal{O}(n^{3})$.
Although the matrix multiplication with large $n$ can still be computationally demanding, the complexity of interval-based inference is substantially lower and remains constant over time, in sharp contrast to the exponentially growing complexity of CCG-based inference.
3.3 Properties of the Interval Inference Attack
The inference attack exhibits certain properties. For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the disturbance and the observation set, as stated below.
Lemma 8.
For any $k\geqslant 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
$p_{k|k}^{y}\leqslant$ $\left(\left|A_{3}\right|+\left|A_{4}\right|\left|A_{2}^{-1}\right|+\left|A_{4}\right|\left|A_{2}^{-1}A_{1}\right|\right)\overline{p}^{x}$
$+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p^{w,x}_{k}+\left|B_{2}\right|p^{w,y}_{k},$ (41)
where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ is a matrix where each element is the absolute value of the corresponding element in $A$, i.e., $|A|=\left[|a_{i,j}|\right]$.
Proof 3.4.
See Appendix B.
Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$. Consequently, if $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level would be low, and the adversary retains little uncertainty after performing the inference attack.
Furthermore, by comparing the predicted and posterior uncertainty sets, as given by (5) and (10), the amount of uncertainty reduction can be quantified by $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, where
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k},$
is the difference between $\mathcal{Y}_{k|k-1}$ and $\mathcal{Y}_{k|k}$. Since $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ can be computed with
$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$ (42)
Therefore, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
Theorem 9.
The amount of uncertainty reduction at $k$ is
$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (43)
satisfying
$\mathrm{Vol}$ $\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}\!-\!c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$.
Proof 3.5.
See Appendix C.
According to Theorem 9, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$. Furthermore, with (42) and Theorem 9, we can show the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$ is bounded in the following lemma.
Lemma 10.
The privacy level can be bounded with the following inequalities,
$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
$\leqslant\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
As a result, we can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$. Moreover, when the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$. Based on this observation, $\mathcal{M}^{x}_{k|k}$ can also be designed to hinder accurate central estimate updates, further improving the privacy level.
4 Privacy Filter Design Problem Using the Volumetric Privacy measure
As discussed previously, an inappropriate choice of the observation set would cause privacy leakage of the private state through inference attacks. To mitigate this risk, we address the privacy filter design problem in this section. The proposed filter determines an appropriate observation set that achieves a desirable balance between preserving the data utility of the public state and ensuring the privacy protection of the private state.
4.1 The Structure of Privacy Filter
We begin by defining the decision domain of the privacy filter as follows. At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the inference set of $X_{k}$ can be computed via,
$\mathcal{Y}_{k|k-1}=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y},$
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$. To maintain the high data utility, the observation set has to satisfy the following constraint,
$\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}.$
To reduce privacy leakage while preserving data utility, we design the privacy filter illustrated in Fig. 2. The design consists of two steps: (1) randomly generate a set $\mathcal{S}_{k|k}^{x}$ such that $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{X}_{k|k-1}$ and $\mathrm{Vol}(\mathcal{S}_{k|k}^{x})\leq\epsilon^{x}$; (2) optimize the observation set $\mathcal{M}^{x}_{k|k}$, which contains $\mathcal{S}_{k|k}^{x}$, to maximize the privacy level. Specifically, in the optimization step, given $\mathcal{S}_{k|k}^{x}$, we maximize the privacy level under the inference attack (6)-(12) by solving
$\mathbf{P_{1}}:$~ $\max_{\mathcal{M}^{x}_{k|k}}~\mathrm{Vol}\!\left(\mathcal{Y}_{k|k}\right)$ (44)
[…]
Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$. In practice, it can be sufficiently small; for instance, it may contain only the true public state $x_{k}$. Consequently, recovering $\mathcal{S}^{x}_{k|k}$ could lead to potential privacy leakage and should therefore be avoided. In the following, we show that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$ due to the randomization operation.
Proposition 11.
The attacker cannot obtain the smaller set $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$.
Proof 4.1.
First, $\mathcal{S}_{k|k}^{x}$ is selected as a random subset of $\mathcal{X}_{k|k-1}$ that contains the true state $x_{k}$. Consequently, $x_{k}$ may reside on the boundary of $\mathcal{S}_{k|k}^{x}$. Next, let $\mathcal{M}_{k|k}^{x,\star}$ denote the optimal observation set. In some cases, $\mathcal{S}_{k|k}^{x}$ may coincide with $\mathcal{M}_{k|k}^{x,\star}$, in which case $x_{k}$ may also lie on the boundary of $\mathcal{M}_{k|k}^{x,\star}$. Therefore, an attacker cannot reconstruct a strictly smaller feasible set containing $x_{k}$ by inverting the optimization process.
As a result, the proposed privacy filter exhibits the following properties: (1) In the absence of inference attacks, the filter output $\mathcal{M}^{x}_{k|k}$ satisfies the utility constraint. (2) In the presence of the inference attack described in (6)–(12), the filter output maximizes the privacy level. (3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$.
Moreover, since the computational complexity of both the inference attack in (6)–(12) and the volume computation increases with the complexity of the set representations, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost. While more sophisticated set representations could improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec. 3.2.
Figure 2: Structure of the proposed privacy filter.
4.2 Randomization
We consider the following random set
${\mathcal{S}}_{k|k}^{x}=\left[\begin{array}[]{c}x_{k}-{\alpha}_{k}\left(x_{k}-\underline{X}_{k|k-1}\right)\\ x_{k}+{\beta}_{k}\left(\overline{X}_{k|k-1}-x_{k}\right)\\ \end{array}\right],$ (52)
where $\alpha_{k}$ and $\beta_{k}$ are uniform random variables with
${\alpha}_{k}\!\in\!\left[\!0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right]\!\!,{\beta}_{k}\!\in\!\left[\!0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right]\!\!.$
Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius from the actual public state $X_{k}=x_{k}$ to the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius from $x_{k}$ to the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ becomes a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state. Also, we can shown $\mathcal{S}^{x}_{k|k}$ satisfies the utility constraint as follows,
$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
$\leqslant$ $\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}$
$+\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
$=$ $\epsilon^{x}.$
We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to improve the privacy level.
4.3 Privacy Filter Optimization
In this subsetion, we demonstrate that the optimization problem $\mathbf{P_{1}}$ based on the interval inference can be solved via linear programming.
Theorem 12.
The privacy filter optimization problem $\mathbf{P_{1}}$ is equivalent to the following linear programming
$\mathbf{P_{2}}:$ $\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$
[…]
$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right$., (63)
where $p_{k-1|k}^{\Delta x}\!\in\mathcal{R}^{n_{x}}$, $Z=X,Y$, $\epsilon^{y}\geqslant 0$ and $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$.
Proof 4.2.
See Appendix D
Figure 3: Inference attack results after applying the optimal privacy filter: (a), (b) Estimated $X_{k}$ and (e), (f) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.01$; (c), (d) Estimated $X_{k}$ and (g), (h) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.5$.
Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3.
5 Numerical Verification
In this section, we study the performance of privacy filter for the production-inventory problem with the following parameters
$A_{1}=\left[\begin{matrix}1.00&0.00\\ 0.00&1.00\\ \end{matrix}\right],A_{2}=\left[\begin{matrix}0.40&0.80\\ 0.60&0.20\\ \end{matrix}\right],$
$A_{3}=\left[\begin{matrix}0.50&-0.90\\ -0.10&-0.10\\ \end{matrix}\right],A_{4}=\left[\begin{matrix}-0.10&-0.90\\ 0.10&0.00\\ \end{matrix}\right],$
$B_{1}=\left[\begin{matrix}-1.00&0.00\\ 0.00&-1.00\\ \end{matrix}\right],B_{2}=\left[\begin{matrix}4.20&0.00\\ 0.00&2.40\\ \end{matrix}\right],$
$\left(\!\mathcal{W}_{k}^{x}\!\right)^{\top}\!\!=\!\left[\!\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\!\right]\!\!,\left(\!\mathcal{W}_{k}^{y}\!\right)^{\top}\!\!=\!\left[\!\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\!\right]\!\!.$
The initial state sets are assumed to be
$\left(\mathcal{X}_{0|-1}\right)^{\top}$ $=\left[\begin{matrix}1.00&0.24&1.20&0.40\\ \end{matrix}\right],$ (64)
$\left(\mathcal{Y}_{0|-1}\right)^{\top}$ $=\left[\begin{matrix}2.40&0.60&3.70&1.30\\ \end{matrix}\right].$ (65)
In our simulation, the initial states are uniformly sampled from the bounded sets (64)-(65). To simulate the approximate periodic fluctuations in demand and productivity, the actual disturbance are set to be
$\left(W_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.88+0.03\cos\left(\frac{2\pi k}{30+7\rho_{k}}\right)&1.94\\ \end{matrix}\right],$
$\!\!\left(W_{k}^{y}\right)\!^{\top}\!$ $=\!\left[\!\begin{matrix}0.944\!+\!0.006\cos\left(\!\frac{2\pi k}{7+2\gamma_{k}}\!\right)\!\!&\!\!0.33+0.094\sin\left(\!\frac{2\pi k}{7+4\tau_{k}}\!\right)\\ \end{matrix}\!\right]\!,$
where $\rho_{k}$, $\gamma_{k}$ and $\tau_{k}$ are uniform random variables in $\left[0,1\right]$. As discussed in Section 2, the production rate is private but the inventory information has to be released.
We first plot the trajectories of system states and their interval tubes in Fig.3 under the optimal privacy filter design for different values of $\epsilon^{x}$. We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$. The pink areas in these figures represent the uncertainty sets of system states. As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate. However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s central estimation of the production rate to become less accurate. This observation numerically verifies Theorem 9, confirming that a higher privacy level prevents the adversary from refining its incorrect central estimation. Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
Figure 4: The truncated Gaussian mechanism.
We also evaluate the utility–privacy trade-off achieved by the proposed optimal privacy-filtering policy and compare it with two benchmark mechanisms: the noiseless quantization method presented in [21] and the truncated Gaussian mechanism for differential privacy introduced in [31]. In the quantization-based approach, the state $x_{k}$ is processed through a static quantizer that satisfies the utility constraint, and the quantization bin containing $x_{k}$ is publicly released as $\mathcal{M}_{k|k}^{x}$. In contrast, the truncated Gaussian mechanism, illustrated in Fig. 4, perturbs the original state $x_{k}$ with additive noise $v_{k}$ drawn from a zero-mean truncated Gaussian distribution supported on the interval $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$ and having variance $\left(\epsilon^{x}\right)^{2}$. The perturbed observation set is then released in the form
$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\frac{\epsilon^{x}}{2n_{x}}\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\frac{\epsilon^{x}}{2n_{x}}\cdot\mathbf{1}_{n_{x}}\end{array}\right],$
where $n_{x}$ denotes the dimension of $x_{k}$ and $\mathbf{1}_{n_{x}}$ is the all-ones vector of length $n_{x}$. Because the additive noise $v_{k}$ is bounded within $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$, the publicly released state is guaranteed to lie within the interval $\mathcal{M}_{k|k}^{x}$.
The privacy-utility trade-off was evaluated by plotting the average privacy level of the production rate against the average utility of the inventory over 100 random trajectories with horizon 100, as illustrated in Fig. 5. For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and its same scaling parameters are applied to the other two mechanisms. The results demonstrate that an increase in inventory utility corresponds to a reduction in the privacy level of the production rate, thereby confirming the intrinsic trade-off between data utility and privacy protection. Besides, the quantization mechanism and the truncated Gaussian mechanism have similar performance in privacy-utility trade-off. Note that these two mechanisms assume static domains of states, while the sets of states in systems subject to disturbance can be estimated and described more precise as addressed by the inference attack. The proposed volumetric approach estimates the time-varying private state set and adjusts the output accordingly.
Furthermore, as discussed in [31], if the volume (e.g., interval length) of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy. However, given the volume constraint, the shape of the public state set can be arbitrary, and certain shapes may lead to substantial volumetric leakage of the private state after inference attacks. The proposed volumetric method explicitly accounts for this by considering the geometry of the set based on the assumed inference attack, not only its volume. Therefore, it achieves higher privacy levels while maintaining lower data distortion compared with the two static mechanisms.
It is worth noting that an adversary could employ more sophisticated estimation techniques, e.g., CCG-based approximation, to infer the private set more accurately from the interval observations provided by the privacy filter. Nevertheless, as shown in Fig. 6, the proposed privacy filter still outperforms the other two mechanisms, leveraging knowledge of the underlying state evolution to reduce conservativeness.
Figure 5: Interval-based inference given the interval privacy filter: the privacy level of the private state and utility of the public state.
Figure 6: CCG-based inference given the interval privacy filter: the privacy level of the private state and utility of the public state.
6 Conclusion
In this paper, we develop a volumetric framework for privacy analysis and defense in dynamic systems subject to bounded disturbance. An inference attack, whereby an adversary estimates the private state, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level. We develop computational methods based on interval analysis, and establish the theoretical properties of the measure. Furthermore, we propose an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method is demonstrated through a production-inventory case study.
It is noted that the performance of the volumetric privacy measure inherently depends on the selected set-membership estimation techniques, and its evaluation accuracy varies with different set representations. Future research will focus on developing approximation methods that ensure improved accuracy, robustness and broader applicability.
Appendix A Proof of Lemma 6
At the time step $k=0$, the adversary only has prior knowledge, i.e., $\mathcal{Y}_{0|-1}$, therefore, its inference set is $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$. Also, since at the time step $k=0$, the backward calibration (6) and (7) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (40).
To prove Lemma 6 for $k\geqslant 1$, we need the following lemma that computes the tightest interval by forward reachability analysis.
Lemma 13.
[32, 33] Consider the static system $S=AM+BW$, where $M$ and $W$ are bounded intervals, the tightest interval for $S$, i.e., $\mathcal{S}$ can be computed as
$\mathcal{S}=\Psi\left(A\right)\mathcal{M}\oplus\Psi\left(B\right)\mathcal{W}.$
Also, we can compute its radius and center via
$p^{s}=\left|A\right|p^{m}$ $+\left|B\right|p^{w},$
$c^{s}=Ac^{m}$ $+Bc^{w}.$
According to Lemma 13, the tightest intervals for (6) and (7) are (24) and (25). Then, the intersection of different intervals, i.e., (8) and (9), can be computed with (28) and (31). Finally, the one-step forward reachable set (10) can be approximated with the tightest interval (36) based on Lemma 13, and the calibrated uncertainty set $\mathcal{X}_{k|k}$ and the tightest prior inference set $\mathcal{Y}_{k|k-1}$ can be approximated similarly.
Appendix B Proof of Lemma 8
According to Lemma 13, the radius of $\mathcal{M}^{y}_{k-1|k}$ can be computed as
$p_{k-1|k}^{m,y}\!=\!\left|\!A_{2}^{-1}\!\right|p_{k|k}^{m,x}\!+\!\left|\!A_{2}^{\!-\!1}\!A_{1}\!\right|p_{k\!-\!1|k-\!1}^{x}\!+\!\left|\!A_{2}^{\!-\!1}\!B_{1}\!\right|p_{k}^{w,x}\!,\!\!$ (66)
where $p^{m,x}_{k|k}$, $p_{k-1|k-1}^{x}$ and $p^{w,x}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{X}_{k-1|k-1}$ and $\mathcal{W}_{k}^{x}$, respectively. Since $\mathcal{Y}_{k-1|k}$ is the intersection result from $\mathcal{M}^{y}_{k-1|k}$ and $\mathcal{Y}_{k-1|k-1}$, the radius of $\mathcal{Y}_{k-1|k}$ is smaller than the radius of $\mathcal{M}^{y}_{k-1|k}$, i.e., $p_{k-1|k}^{y}\leqslant p_{k-1|k}^{m,y}$. Also, the radius of $\mathcal{Y}_{k|k}$ can be computed as
$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$ (67)
By substituting (66) and $p_{k-1|k}^{y}\leqslant p_{k-1|k}^{m,y}$ into (67), we have
$p_{k|k}^{y}\leqslant\left|A_{3}\right|p_{k-1|k}^{x}+\left|B_{2}\right|p_{k}^{w,y}+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x}$
$+\left|A_{4}\right|\left(\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}\right).$
Since $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$ and $\mathcal{X}_{k-1|k}$ is a subset of $\mathcal{M}_{k-1|k-1}^{x}$, we have $p_{k-1|k}^{x}\leqslant p^{m,x}_{k-1|k-1}\leqslant\overline{p}^{x}$ for any $k\geqslant 1$, thus we have (8).
Appendix C Proof of Theorem 9
The difference set $\Delta\mathcal{Y}_{k|k}$ is computed as,
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}\oplus\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
where
$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}\setminus\mathcal{X}_{k-1|k}$
$=\left[\begin{array}[]{c}\min\left\{\underline{X}_{k-1|k-1}-\underline{M}_{k-1|k}^{x},0\right\}\\ \max\left\{\overline{X}_{k-1|k-1}-\overline{M}_{k-1|k}^{x},0\right\}\\ \end{array}\right],$ (70)
$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}\setminus\mathcal{Y}_{k-1|k}$
$=\left[\begin{array}[]{c}\min\left\{\underline{Y}_{k-1|k-1}-\underline{M}_{k-1|k}^{y},0\right\}\\ \max\left\{\overline{Y}_{k-1|k-1}-\overline{M}_{k-1|k}^{y},0\right\}\\ \end{array}\right].$ (73)
Therefore, the volume of the difference set is (43).
With Lemma 13, we have
$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$
where the radius $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ can be computed via
$2p_{k-1|k}^{\Delta z}$
$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}-\min\left\{\underline{Z}_{k-1|k-1}-\underline{M}_{k-1|k}^{z},0\right\}$
$=$ $\max\left\{0,\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}+\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},\right$.
$\!\!\!\!\overline{Z}_{k-\!1|k-\!1}\!-\!\overline{M}_{k-\!1|k}^{z},\underline{M}_{k-1|k}^{z}\!-\!\underline{Z}_{k-\!1|k-\!1}\left.\right\},\text{for}\,Z=X,Y,\!\!$ (74)
which satisfies $p_{k-1|k}^{\Delta z}\geqslant 0$. As a result, we have
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
$\overset{(a)}{\leqslant}\left\|A_{3}\right\|_{1}\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|_{1}\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
$=\left\|A_{3}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
where $(a)$ is due to $p_{k-1|k}^{\Delta x}\geqslant 0$ and $p_{k-1|k}^{\Delta y}\geqslant 0$.
Besides, given an interval $\mathcal{X}$, we can express it with its center point and radius, i.e., $\mathcal{X}=\left[\begin{array}[]{c}\frac{c-p}{2}\\ \frac{c+p}{2}\end{array}\right]$. Therefore, we have
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
$\geqslant$ $\left\|\overline{Y}_{k|k}+\underline{Y}_{k|k}-\left(\overline{Y}_{k|k-1}+\underline{Y}_{k|k-1}\right)\right\|_{1}$
$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
Appendix D Proof of Theorem 12
To maximize the privacy level, it is equivalent to minimize the amount of uncertainty reduction since we have $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, where the prior uncertainty set $\mathcal{Y}_{k|k-1}$ is fixed at time step $k$.
Besides, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=\left\|p_{k|k}^{\Delta y}\right\|_{1}=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$, where the elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ are non-negative vectors as shown in (C). Therefore, we can replace the objective function with the slack variable $\epsilon^{y}$ and add $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\leqslant\epsilon^{y}$ as a new constraint, and then minimize $\epsilon^{y}$.
Since $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (28) and (31) with the constraints of difference sets (C). Also, the objective function increases with any elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ since the elements of $p_{k-1|k}^{\Delta x}$, $p_{k-1|k}^{\Delta y}$, $\left|A_{3}\right|$ and $\left|A_{4}\right|$ are non-negative. As a result, we can replace the constraint of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, i.e., (C), with inequalities (63), and let $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ be decision variables.
Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.
Finally, the objective and the constraints are linear functions of the decision variables, thus, the optimal privacy filter can be obtained by solving the linear programming $\mathbf{P_{2}}$.